OpenStudy (anonymous):

a women pulls a 7.87 kg suitcase. initially at rest with a 29.2 N force along the handle at 62.7 degress. friction pulls back at 11.7 N and the suitcase moves 8.44 m what is the total work done?

OpenStudy (turingtest):

\[\vec F=\vec F_{push}-\vec F_{friction}\]and work is\[\vec F\cdot \vec d=Fd\cos\theta\]

OpenStudy (turingtest):

where \(\vec F\) is the total force

OpenStudy (anonymous):

cosine is 0?

OpenStudy (turingtest):

what is the angle \(\theta\) between \(\vec F\) and \(\vec d\) ?

OpenStudy (anonymous):

147.7

OpenStudy (turingtest):

oh I'm sorry, I misunderstood the question starting over :p what is the formula for friction force?

OpenStudy (anonymous):

mgh+1/2mv^2=mgh+1/2mv^2 thats the equations it give us

OpenStudy (turingtest):

sounds like you're looking up equations in your notes and books without really knowing what they mean what you wrote there is conservation of energy for a falling object, and that obviously is not what we need right now. What I said earlier is true: total forge times distance is work done. In order to know that you need to know the force of friction. Again, what is the formula for the force of friction?

OpenStudy (turingtest):

total force*

OpenStudy (anonymous):

it doesnt tell me

OpenStudy (turingtest):

are you sure that the problem is worded exactly correctly? is it being pulled 60 degrees up from the horizontal, or down, or side to side...?

OpenStudy (turingtest):

or is there a picture maybe? because the question is a bit incomplete

OpenStudy (anonymous):

no its just the question

OpenStudy (turingtest):

well I assume that they are pulling 62.7 deg up, in which case you need to find the normal force of the object against the ground, from which you can find the friction force. can you do that?